# CoViS-Net: A Cooperative Visual Spatial Foundation Model for Multi-Robot Applications

### v1

@@ IV-C Training
We use Pytorch Lightning to train our model for 15 epochs on two NVidia A100 GPUs. This training takes approximately 24 hours. We optimize the weights using the AdamW optimizer and use the 1cycle learning rate annealing schedule throughout the training and configured weight decay. We choose the best model based on the performance on the validation set. We provide further details about the training procedure in the supplementary materials.
After training the model on $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$, we freeze the weights and perform no further sim-to-real finetuning, thus zero-shot transferring from simulation to real-world evaluation.
V Experiments and Results
We demonstrate the performance of our trained models on a simulated testset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ and new real-world test sets $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$, introduced in this section. First, we outline the metrics we use. Then, we discuss our simulation experiments and results. Next, we collect real-world images using our multirobot system, and analyze the model performance. Finally, we compile and deploy our model on a real-time trajectory tracking task, and report our findings.
@@ V Experiments and Results
Metrics: We compute the absolute Euclidean distance between the predicted and ground truth poses and the geodesic distance between rotations as
@@
$D_{\mathrm{rot}}(\mathbf{R}_{i,ij},\hat{\mathbf{R}}_{i,ij})$ $=4\cdot\mathrm{arcsin}\left(\frac{1}{2}d_{\text{quat}}\left(\hat{\mathbf{R}}_{i,ij},\mathbf{R}_{i,ij}\right)\right)$
@@
For simulation results, we also report the Dice as well as the Intersection over Union (IoU) metric.
V-A Simulation
We conducted a series of experiments across various models trained with different parameters to showcase our model’s efficiency. The quantitative results are presented in Tab. I and Tab. II, with qualitative findings in Fig. 2.
TABLE I: Ablation study over the number of patches and size of features for each patch. We report the BEV performance as well as the median error for poses on the dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$. The size of the embedding $\mathbf{E}_{i}$ is reported in bytes.
Model BEV Median Pose Err.
S F Mem Dice IoU
[…]
Tab. I details the performance across seven different experiments on the simulation test set $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$, as outlined in section Sec. IV-A. We evaluated the impact of image embedding sizes $\mathbf{E}_{i}$ on communication efficiency, reporting sizes based on 16-bit floats. The Intersection over Union (IoU) and Dice metrics for the BEV representation of the local environment, alongside median pose estimates for all samples in the dataset, are provided. It’s important to note the limitation in reporting estimation feasibility between nodes due to Habitat’s lack of specific information (e.g., samples facing the same direction might be in adjacent rooms, unseen by the dataset).
Training one model with $F=48$ and $S=128$ across $N=5$ different seeds revealed a very low average standard deviation of $0.011374$, leading us to report results for only $N=1$ seeds in the table for clarity.
The experiments demonstrate that the sequence length $S$, acting as the context size for each image, significantly influences BEV prediction and pose estimate performance more than feature size $F$. We set $S$ to 1 for two experiments: (1) we increase $F$ to match the size of the embeddings for the other experiments, and (2) we decrease $F$, which effectively reduces the transformer to a CNN-only architecture by removing the attention mechanism. This resulted in a marked decrease in performance for both BEV and pose estimations, underscoring the critical role of attention in our model. The best-performing model with $S=256$ achieved an BEV IoU of $0.571$, a Dice score of $0.691$, and median pose accuracy of $36$ cm and $8.3^{\circ}$. Conversely, models with $S=1$ showed significant degradation to near unusable BEV predictions and pose errors on the scale of meters. Models with $F=24$ and $S=64$ exhibited slightly inferior performance compared to those with larger $F$ or $S$.
TABLE II: Ablation over different modes for the BEV prediction on the simulation testset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
@@ V-A Simulation
Lastly, we assess the effectiveness of BEV predictions, focusing on the contribution of pose predictions to enhancing BEV accuracy, in an ablation study detailed in Tab. II. We conduct three experiments: (1) A model trained without incorporating any pose prediction into the BEV aggregator. This model can estimate local information and, to a degree, information aggregated through neighbors. However, it must implicitly learn to predict the relative poses between observations to successfully merge the BEV representations from multiple agents. (2) A model trained as described in this work, integrating pose estimates with BEV predictions and aggregating this information from neighbor nodes. For this experiment, we use the model configuration with $F=48$ and $S=128$. (3) A BEV model trained with ground truth pose information from the simulation training set $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$. This model’s performance is considered an upper limit, assuming perfect pose estimation. Evaluation on the simulation test set $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ indicates our model performs between the two baselines. It significantly outperforms the pose-free model with an $8.75\%$ improvement in accuracy, while the model using ground truth poses shows an $18.31\%$ performance increase compared to the baseline without pose information. These results highlight the substantial benefit of incorporating pose predictions into BEV estimation, positioning our approach as a notably effective middle ground between the two extremes.
V-B Real-World
In this section, we detail our robot setup, custom real-world test set, and model evaluation on this dataset. We run a series of experiments with quantitative results in Tab. III, Tab. VI, Tab. IV, Tab. V, and Fig. 5 and qualitative findings in Fig. 4, Fig. 7, and Fig. 6.
Robot Platform: Our setup includes four DJI RoboMaster S1 robots, each equipped with an NVidia Jetson Orin NX 16GB and a forward-facing calibrated Raspberry Pi HQ camera. The camera’s raw field of view is approximately $170^{\circ}$, rectified to $120^{\circ}$ using the OpenCV Fisheye module, with an image resolution of $224\times 224$ px.
Real-World Testset Collection: Our custom real-world test set $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ was collected in eight office-like indoor scenes, tailored to our robot platform’s requirements. Unlike existing datasets like Seven Scenes, which have a constrained FOV of $53.8^{\circ}$, our dataset is designed for our model’s application, reflecting dynamic environmental changes through multiple agents. It represents one of the first multi-agent datasets collected with locally connected devices.
Data collection involved a base station with dual cameras and a 2D Lidar, plus a distinctively shaped marker on each robot. We manually navigated robots around the base station to capture a variety of relative poses, then used a custom tool to label Lidar data for pose extraction. In total, we collect 5692 labeled samples. To balance out the distribution of relative positions in the dataset, we sample image pairs to ensure a uniform distribution of relative positions in the range of 0 m to 2 m. We sample a similar number of pairs from all scenes, resulting in a total of 14008 sets of samples each with three images, resulting in a total of $14008\cdot 2\cdot 3=84048$ pose edges. $32\%$ of these ground truth edges have no visual overlap. We show a sample of each scene from the dataset in Fig. 4. Our dataset covers a diverse and challenging range of indoor scenes, as well as one outdoor scene. We include more details about the testset in the supplementary material.
(a) Corridor A
(b) Corridor B
(c) Office A
(d) Office B
(e) Study A
(f) Study B
(g) Outdoor
(h) Sunny
Figure 4: We collect the real-world dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ from five unique scenes in an indoor office building. The scenes show different challenges, from cluttered environments over scenes with high ceilings to sunny floors.
Metrics: In Tab. III, Tab. VI, Tab. IV and Tab. V, we categorize the median pose error into the categories Invisible, Visible and Invisible Filtered. Visible and Invisible edges are determined through the FOV overlap (an edge is considered invisible if the angle $D_{\mathrm{rot}}(\mathbf{R}_{i,ij},\mathbf{0})>\mathrm{FOV}$). Invisible Filtered samples use the predicted position uncertainty $f_{\textrm{p}}^{\sigma}{ij}$ of our model to reject pose estimations with high uncertainty based on a threshold computed with the Youden’s index [66].
Results: Evaluation on the real-world dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ focuses on position and rotation accuracy, excluding quantitative BEV performance due to unavailable ground truth data.
We first report results similar to Tab. I on $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ in Tab. III. The results are in line with the results reported for $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$, with a distinct correlation between the number of patches $S$ and features $F$ and the performance of pose estimation. Notably, the performance in estimating rotations does not significantly vary between visible and invisible samples, demonstrating our model’s capability to robustly estimate poses without necessitating FOV overlap. The model exhibiting the highest accuracy achieves a median localization error of 97 cm and $7.9^{\circ}$ for invisible samples, and 33 cm and $5.8^{\circ}$ for visible samples. Note that the samples in the dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ are at most 2 m apart, but the pose prediction is three-dimensional, resulting in a worst-case error of up to 4 m.
TABLE III: Ablation study over the number of patches and size of features for each patch. We report the BEV performance as well as the median error for poses on the dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$. The size of the embedding $\mathbf{E}_{i}$ is reported in bytes.
S F Mem Invisible Filtered Invisible Visible
[…]
In all further evaluations, we detail the outcomes for the model configuration with $S=128$ and $F=48$. This model was selected for its balanced trade-off between BEV and pose estimation accuracy and manageable embedding size.
@@ V-B Real-World
These findings demonstrate our model’s capability to reliably estimate poses, even with minimal overlap, and to accurately predict higher variances under such conditions. We attribute this performance to the robust representational ability of the DinoV2 encoder and vision transformer models at large. The model appears to go beyond mere feature matching, instead achieving a sophisticated scene understanding that allows it to “imagine” unseen scene portions based on learned data priors, thereby facilitating pose estimation.
In Tab. IV, we investigate the performance of our model across the different scenes in the testing dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$. We provide a detailed breakdown of the median pose error for Invisible, Invisible Filtered and Visible cases for each of the eight scenes. While the pose error for visible samples is consistently between 22 cm/5.2°and 48 cm/7.5°, there are more outliers for invisible samples. This is due to some scenes having more or less symmetry, thus making it more challenging for the model to accurately predict poses. The pose prediction for the outdoor scene are least accurate, which is to be expected since the model was trained on indoor scenes. The Sunny scene has severe shadows and direct sunlight, while still providing low pose errors. The pose predictions are useful in all scenarios, demonstrating impressive generalization ability across a wide range of scenes.
TABLE IV: The median pose estimation error of our model on each of the eight $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ scenes.
@@
Next, we evaluate the magnitude of pose prediction errors over different thresholds of distances between image pairs in Tab. V. We consider five bands of distances separated in 0.4 m chunks, ranging up to 2m. Note that the worst-case pose error is $2\cdot d$ due to the three-dimensional pose predictions. While the error increases with an increasing distance between samples, the rotational error in particular stays low, and the position errors stay useful, ranging from 12 cm to 65 cm for visible samples, 34 cm to 134 cm for invisible filtered samples (samples that are not visible but for which he model predicts a high confidence) and 76 cm to 153 cm for all invisible samples.
TABLE V: We show median pose prediction errors for different distance thresholds $d$ on the dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$.
@@
We provide a qualitative sample in Fig. 6, where additionally to the simulation sample in Fig. 2, we also visualize the prediction uncertainty. More samples can be found in the supplementary material.
@@
Figure 6: By leveraging priors, such as the way humans lay out their rooms, our model can still provide rough relative position even without overlapping fields of view. This is generally impossible to do using classical approaches. We visualize a sample from our custom real-world testset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ for three nodes, similar to the simulation sample in Fig. 2. We do not show the ground-truth BEV segmentation since we do not have labels for this. Additionally to the pose predictions $\hat{\mathbf{p}}_{i,ij},\hat{\mathbf{R}}_{i,ij}$, we also visualize the predicted uncertainties $\sigma^{2}_{i}{ij}$.
Finally, we conducted a baseline comparison with two feature-matching-based approaches, as summarized in Table VI. Both baseline methods extract and match features across two images to estimate the essential matrix, from which a rotation and a scale-less translation vector are derived. The first baseline employs OpenCV for ORB feature extraction followed by a brute-force matching technique. The second baseline, LightGlue, utilizes a neural network-based approach for feature extraction and matching. Given that both baselines estimate scale-less transformations, we adopted the same metric as used in LightGlue [23]. This metric computes the AUC at thresholds of 20, 45, and 90 degrees based on the maximum error in rotation and translation. Our findings indicate that our model surpasses both baselines in terms of performance across image pairs with and without visual overlap. It is important to note that while both baselines necessitate visual overlap to function effectively, our approach does not require such overlap, highlighting its robustness and applicability in a broader range of scenarios.
TABLE VI: We report the AUC metric at 20, 45 and 90°on a first-principles OpenCV and an NN-based baseline. We report the AUC metric at 20, 45 and 90 degree.
Invisible Visible
AUC@ 20 45 90 20 45 90
ORB/OpenCV [42] 0.16 1.60 7.43 9.41 21.92 38.03
LightGlue [23] 0.02 0.09 0.31 23.94 39.82 55.14
Ours 9.22 24.40 45.74 32.74 58.53 76.01
V-C Real-World Pose Control
In this section, we describe our real-world (online) multi-robot control experiment setup, including model deployment, wireless communication strategies, controllers, and results reported in Fig. 7, Fig. 8, Tab. VIII and Sec. V-C.
@@ V-C Real-World Pose Control
0s1s2s3s4s5s
[…]
TABLE VII: We report the mean and absolute tracking error for both leaders for all three trajectories, as well as average velocities. The errors are consistent with the results reported in Tab. III.
Trajectory Robot Mean Abs. Median Vel
Fig. 8 dyn. A $38$ cm, $5.6^{\circ}$ $38$ cm, $4.8^{\circ}$ 0.59 m/s
B $28$ cm, $5.1^{\circ}$ $25$ cm, $4.7^{\circ}$ 0.58 m/s
Fig. 8 stat. A $51$ cm, $5.2^{\circ}$ $48$ cm, $5.3^{\circ}$ 0.60 m/s
B $28$ cm, $5.6^{\circ}$ $26$ cm, $5.4^{\circ}$ 0.61 m/s
Rect. dyn. A $41$ cm, $4.4^{\circ}$ $42$ cm, $4.4^{\circ}$ 0.32 m/s
B $29$ cm, $5.1^{\circ}$ $27$ cm, $5.1^{\circ}$ 0.81 m/s
Figure 8: We evaluate the tracking performance of our model and uncertainty-aware controller on three reference trajectories, with two follower robots (in blue and orange), positioned left and right of the leader robot (in green). For each trajectory, we report the tracking performance over time for position and rotation, as well as the predicted uncertainties. The left trajectory (1) is a Figure Eight with the leader always facing in the same direction. The middle trajectory (2) is similar to the first, but the leader is always facing the direction it is moving. In this trajectory, Follower A starts facing the opposite direction of the leader, resulting in high position variance and low rotation variance, which decreases as the robot rotates, facing the same direction as the leader. A spurious erroneous rotation prediction is accompanied by a high uncertainty. Note that as the leader moves through the trajectory, there is always one inside and one outside follower robot. This role changes depending on the section of the figure, resulting in these follower trajectories. The right trajectory (3) is a rounded rectangle. We show the trajectory for 120 s and the tracking error for the first 60 s. The tracking performance is consistent across multiple runs.
Model deployment: We deploy a GNN-based model, selecting $F=24$ and $S=128$ for its balance between embedding size and performance, as reported in Tab. III. After training centralized in PyTorch, we compile $f_{\textrm{enc}}$, $f_{\textrm{pose}}$, and $f_{\textrm{agg}}$ for decentralized deployment. Using TensorRT, a proprietary NVidia tool for ahead-of-time compilation of neural networks, we achieve sub-50 ms processing times. We benchmark the runtime performance of the model on the Jetson Orin NX in Tab. VIII by averaging 100 sequential forward passes. We compare 16-bit and 32-bit float evaluated on the GPU or CPU and compilation with TensorRT. The model $f_{\textrm{enc}}$, generating the communicated embedding $\mathbf{E}_{i}$, has a total of 30 M parameters, and takes ca. 20 ms for one forward pass as FP16 optimized with Torch TensorRT on the GPU. This is 50×50$\times$50 × faster than running the same model on the CPU. We also note that changing from 32-bit floats to 16-bit floats results in a 2.75×2.75$\times$2.75 × speedup. The model $f_{\textrm{pose}}$, processing a pair of embeddings into a relative pose, has a total of 6 M parameters and takes ca. 4 ms for one forward pass. The model is run in a custom C++ environment integrated with ROS2 Humble and an ad-hoc WiFi network for inter-node communication to communicate image embeddings $\mathbf{E}_{i}$, achieving a 15 Hz processing rate for the whole pipeline, including image pre-processing, model evaluation, communication and pose prediction.
TABLE VIII: Model runtime evaluation for $S=128$ and $F=24$.
Type Mode Dev. $f_{\textrm{enc}}$ $f_{\textrm{pose}}$
FP16 TRT GPU 19.22 ms $\pm$0.92 ms 4.15 ms $\pm$0.49 ms
FP16 JIT GPU 36.79 ms $\pm$1.33 ms 8.04 ms $\pm$0.73 ms
FP32 JIT GPU 99.32 ms $\pm$1.56 ms 13.57 ms $\pm$1.10 ms
FP32 JIT CPU 952.04 ms $\pm$116.72 ms 248.41 ms $\pm$45.98 ms
Ad-hoc WiFi: Image embeddings are sent between nodes at a rate of 15hz, with each being just over 6KiB. The fully distributed nature of our system matches well with a broadcast communications topology, which theoretically can permit operation at scale. Broadcast messaging has a significant drawback, however, in that modern wireless data networking standards (IEEE 802.11/WiFi) fallback to very low bit rates to ensure maximum reception probability.
A higher level system, such as ROS2 communications middleware package CycloneDDS, can manage retransmissions, however currently available systems are not well suited to the requirements of a decentralized robotic network. In particular, they cannot detect network overload conditions, which can result in retransmissions that further load the network; finally resulting in a rapid increase in packet loss rates which renders the network as a whole unusable. While it is possible to tune such communications packages manually, the best parameter values are sensitive to highly variable deployment configuration parameters including ambient wireless network traffic. Manual tuning of network hardware, such as forcing increased broadcast bit rates, can alleviate this problem but is generally not as resilient at scale.
Our custom messaging protocol includes the ability to dynamically backoff messaging load based upon detected network conditions. Our system is also uses a shared slot TDMA message scheduling system which works with the 802.11 frame scheduler to maximize frame reception probability without compromising overall data rates.
Uncertainty-aware relative pose control: We utilize a PD controller for relative pose control of follower robots, implementing independent controllers for position and orientation. The controllers adjust based on the predicted uncertainty, deactivating if uncertainty exceeds certain thresholds to prevent inaccurate operations. This ensures that even with unreliable pose estimates, the robot can maintain its orientation toward the leader until a reliable estimate becomes available.
Experiment: We conduct qualitative experiments in a wide range of different environments, including outdoor, and quantitative experiments in a large room equipped with a motion capture system for accurate pose tracking. Two follower robots are programmed to maintain a fixed distance of $0.6$ m from a leader robot executing predefined trajectories. For the quantitative experiments, we assess the system’s tracking accuracy by comparing the ground truth and estimated relative poses, focusing on position and rotation errors and the predicted uncertainties for each dimension, using the same metrics introduced in Sec. V. Our model computes one uncertainty per 3D dimension, which we report as mean.
The three trajectories are two Figure Eight trajectory, one for which we keep the orientation of the leader constant, and one for which the leader is always facing in the direction it is moving. The third trajectory is a rounded rectangle, again with the leader facing the direction it is moving.
Results: Fig. 7 shows snapshots from four different real-world deployments, and Fig. 8 shows the trajectories of all robots for all quantitative experiments, as well as a visualization of the position and rotation error over time. We provide additional snapshots from more real-world experiments in the appendix.
The qualitative experiments show a robust system that generalizes to a wide range of scenarios, including outdoors, albeit the scale prediction is less reliable in outdoor settings. Our model was trained exclusively on indoor data and generalizes to outdoor data due to the wide range of training data used on the DinoV2 encoder. The performance can be significantly improved by training on outdoor samples.
The tracking error in the quantitative experiments oscillates throughout the trajectories as the followers reactively adjust their position to the movements of the leader. Additionally, Figure 8 dynamic starts with Follower A facing the opposite direction of the Leader, resulting in higher position error and uncertainty while the rotation error is low. As the Follower rotates, the position uncertainty decreases. During this process, the model generates an erroneous rotation prediction that is accompanied with a spike in the uncertainty prediction. We note that the tracking error for Figure 8 static differs between both agents. This is due to the lack of structure in the lab environment and it affects both robots differently, resulting in better tracking performance for Follower A. The tracking error is consistent for the other trajectories. Sec. V-C reports quantitative evaluations for the same results, which are in line with the results reported in Tab. III. Note that the velocities reported are average velocities of the follower robots. The velocities are consistent for both Figure 8 trajectories, but differ for the rectangle trajectory, as the inner robot moves more slowly than the outer robot.
VI Conclusion
In this work, we introduced a novel visual spatial foundation model for multi-robot control applications, based on an existing computer vision foundation model. From monocular camera images, our model outputs accurate relative pose estimates and BEVs in the local frame by aggregating information from other robots. We validated our model on a custom real-world testset collected in various locations of an indoor office building. Then, we deployed our model onto robots, demonstrating real-time multi-robot control. We find that our model can accurately predict poses, even when there are no pixel-level correspondences, and correctly predict BEVs for obscured regions. In the real-world, we demonstrate accurate trajectory tracking, paving the way for more complex vision-only robotics tasks. Our model is able to estimate relative poses with a median pose error of up to $33$ cm and $5.8^{\circ}$ for visible edges and 97 cm and $5.9^{\circ}$ for invisible edges on our real-world dataset containing $84048$ relative poses. This is roughly equivalent to one body length of our robot. Using these estimated poses, our model increases the BEV prediction score by $8.75\%$ compared to a multi-node aggregation scheme without poses.
In future work, we plan to investigate how to apply this model for other downstream tasks, and in combination with learning-based multi-agent control policies.
Acknowledgments
This work was supported in part by European Research Council (ERC) Project 949949 (gAIa). J. Blumenkamp acknowledges the support of the ‘Studienstiftung des deutschen Volkes’ and an EPSRC tuition fee grant. We also acknowledge a gift from Arm.
Appendix A Training
A-A Dataset
The samples of the datasets $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$ $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ and $\mathcal{D}_{\mathrm{Val}}^{\mathrm{Sim}}$ are randomly sampled to be within orientation range $[-\frac{\pi}{32},\frac{\pi}{32}]$ for roll and pitch, and within a height of 0.1 to 0.3 meters, representing the camera mounted on the robot.
A-B Model architecture
The encoder $f_{\textrm{enc}}$ consists of the pre-trained DinoV2 vision transformer and two more transformer layers previously described as $f_{\textrm{trans}}^{\textrm{enc}}$. The input feature size is $224\times 224$, of which DinoV2 generates an embedding of size $S=256$ and $F=384$. We add a linear layer that transforms $F$ down to the parameter described in this paper (e.g. $24$) while maintaining $S$. The output of this is then cropped to the $S$ described in this paper (e.g. 128) by cropping the sequence length at that position.
The pairwise pose encoder $f_{\textrm{pose}}$ consists of transformer $f_{\textrm{trans}}^{\textrm{pose}}$, which is assembled from five transformer blocks, and four MLPs $f^{\mu}_{\textrm{p}}$, $f_{\textrm{p}}^{\sigma}$, $f^{\mu}_{\textrm{R}}$, $f_{\textrm{R}}^{\sigma}$. We first project the embedding $\mathbf{E}_{i}$ to a larger feature size $F=192$ using a linear layer. Each transformer layer has $F=192$ in and out features. The learnable transformer positional embedding $\mathbf{S}_{\mathrm{pose}}$ is initialized to match $F$ and $S$, with normal distribution initialization of $\sigma=0.02$. The MLPs $f^{\mu}_{\textrm{p}}$, $f_{\textrm{p}}^{\sigma}$, $f^{\mu}_{\textrm{R}}$, $f_{\textrm{R}}^{\sigma}$ are represented by one linear layer mapping the output of the transformer from $F=192$ to $F=17$ (position: 3, position uncertainty: 3, orientation: 10, orientation uncertainty: 1).
A-C Loss function
A-D Training
We show the train and validation loss for $N=5$ training runs with different seeds in Fig. 9.
Appendix B Real-World Dataset
We visualize the five different scenes we collected the dataset from as well as the data distribution from each respective scene in Fig. 12. We explain the setup of our data collection unit in Fig. 10.

### v2

@@ A.4 Training
We use Pytorch Lightning to train our model for 15 epochs on two NVidia A100 GPUs. This training takes approximately 24 hours. We optimize the weights using the AdamW optimizer and use the 1cycle learning rate annealing schedule throughout the training and configured weight decay. We choose the best model based on the performance on the validation set.
We show the train and validation loss for $N=5$ training runs with different seeds in Fig. 8. Training with $F=48$ and $S=128$ across $N=5$ different seeds reveals a low average standard deviation of $0.011374$, leading us to report results for only $N=1$ seeds in Tab. 1.
Appendix B Experiments and Results
B.1 Metrics
@@
We compute the absolute Euclidean distance between the predicted and ground truth poses and the geodesic distance between rotations as
@@
$D_{\mathrm{rot}}(\mathbf{R}_{i,ij},\hat{\mathbf{R}}_{i,ij})$ $=4\cdot\mathrm{arcsin}\left(\frac{1}{2}d_{\text{quat}}(\hat{\mathbf{R}}_{i,ij},\mathbf{R}_{i,ij})\right)$
@@
B.2 Simulation
We provide additional quantitative experiments in Fig. 9 and Tab. 3. We visualize four additional qualitative samples from the simulation dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ in Fig. 10, Fig. 11, Fig. 12, Fig. 13, Fig. 14 and Fig. 15.
Figure 10: Sample A from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 11: Sample B from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 12: Sample C from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 13: Sample D from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 14: Sample E from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
Figure 15: Sample F from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
BEV Experiments: We assess the effectiveness of BEV predictions, focusing on the contribution of pose predictions to enhancing BEV accuracy, in an ablation study detailed in Tab. 3.
Table 3: Ablation over different modes for the BEV prediction on the simulation testset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$.
@@ B.2 Simulation
We conduct three experiments: (1) A model trained without incorporating any pose prediction into the BEV aggregator. This model can estimate local information and, to a degree, information aggregated through neighbors. However, it must implicitly learn to predict the relative poses between observations to successfully merge the BEV representations from multiple agents. (2) A model trained as described in this work, integrating pose estimates with BEV predictions and aggregating this information from neighbor nodes. For this experiment, we use the model configuration with $F=48$ and $S=128$. (3) A BEV model trained with ground truth pose information from the simulation training set $\mathcal{D}_{\mathrm{Train}}^{\mathrm{Sim}}$. This model’s performance is considered an upper limit, assuming perfect pose estimation. Evaluation on the simulation test set $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Sim}}$ indicates our model performs between the two baselines. It significantly outperforms the pose-free model with an $8.75\%$ improvement in accuracy, while the model using ground truth poses shows an $18.31\%$ performance increase compared to the baseline without pose information. These results highlight the substantial benefit of incorporating pose predictions into BEV estimation, positioning our approach as a notably effective middle ground between the two extremes.
B.3 Real-World
We provide additional quantitative experiments in Fig. 16, Tab. 4 and Tab. 5. We visualize four additional qualitative samples from the real-world dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ in Fig. 17 and Fig. 18.
@@ B.3 Real-World
In Tab. 4, we investigate the performance of our model across the different scenes in the testing dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$. We provide a detailed breakdown of the median pose error for Invisible, Invisible Filtered and Visible cases for each of the eight scenes. While the pose error for visible samples is consistently between 22 cm/5.2°and 48 cm/7.5°, there are more outliers for invisible samples. This is due to some scenes having more or less symmetry, thus making it more challenging for the model to accurately predict poses. The pose prediction for the outdoor scene are least accurate, which is to be expected since the model was trained on indoor scenes. The Sunny scene has severe shadows and direct sunlight, while still providing low pose errors. The pose predictions are useful in all scenarios, demonstrating impressive generalization ability across a wide range of scenes.
Table 4: The median pose estimation error of our model on each of the eight $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ scenes.
@@
Next, we evaluate the magnitude of pose prediction errors over different thresholds of distances between image pairs in Tab. 5. We consider five bands of distances separated in 0.4 m chunks, ranging up to 2m. Note that the worst-case pose error is $2\cdot d$ due to the three-dimensional pose predictions. While the error increases with an increasing distance between samples, the rotational error in particular stays low, and the position errors stay useful, ranging from 12 cm to 65 cm for visible samples, 34 cm to 134 cm for invisible filtered samples (samples that are not visible but for which he model predicts a high confidence) and 76 cm to 153 cm for all invisible samples.
Table 5: We show median pose prediction errors for different distance thresholds $d$ on the dataset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$.
@@
Figure 17: We visualize a sample from our custom real-world testset $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$ for three nodes, similar to the simulation sample in Fig. 2. We do not show the ground-truth BEV segmentation since we do not have labels for this. Additionally to the pose predictions $\hat{\mathbf{p}}_{i,ij},\hat{\mathbf{R}}_{i,ij}$, we also visualize the predicted uncertainties $\sigma^{2}_{\mathrm{p},i}{ij}$ and $\sigma^{2}_{\mathrm{R},i}{ij}$.
Figure 18: Sample B and C from $\mathcal{D}_{\mathrm{Test}}^{\mathrm{Real}}$.
B.4 Real-World Pose Control
Ad-hoc WiFi: Image embeddings are sent between nodes at a rate of 15hz, with each being just over 6KiB. The fully distributed nature of our system matches well with a broadcast communications topology, which theoretically can permit operation at scale. Broadcast messaging has a significant drawback however, in that modern wireless data networking standards (IEEE 802.11/WiFi) fallback to very low bit rates to ensure maximum reception probability.
A higher level system, such as ROS2 communications middleware package CycloneDDS can manage retransmissions, however currently available systems are not well suited to the requirements of a decentralized robotic network. In particular they lack the ability to detect network overload conditions, which can result in retransmissions that further load the network; finally resulting in a rapid increase in packet loss rates which renders the network as a whole unusable. While it is possible to tune such communications packages manually, the best parameter values are sensitive to highly variable deployment configuration parameters including ambient wireless network traffic. Manual tuning of network hardware, such as forcing increased broadcast bit rates, can alleviate this problem but are generally not resilient as scale.
Our custom messaging protocol is targeted to this specific use case, and includes the ability to dynamically backoff messaging load based upon detected network conditions. Our system is also uses a shared slot TDMA message scheduling system which works with the 802.11 frame scheduler to maximize frame reception probability without compromising overall data rates.
Results: We benchmark the runtime performance of the model on the Jetson Orin NX in Tab. 6 by averaging 100 sequential forward passes. We compare 16-bit and 32-bit float evaluated on the GPU or CPU and compilation with TensorRT, a proprietary NVidia tool for ahead-of-time compilation of neural networks. The model $f_{\textrm{enc}}$, generating the communicated embedding $\mathbf{E}_{i}$, has a total of 30 M parameters, and takes ca. 20 ms for one forward pass as FP16 optimized with Torch TensorRT on the GPU. This is 50×50$\times$50 × faster than running the same model on the CPU. We also note that changing from 32-bit floats to 16-bit floats results in a 2.75×2.75$\times$2.75 × speedup. The model $f_{\textrm{pose}}$, processing a pair of embeddings into a relative pose, has a total of 6 M parameters and takes ca. 4 ms for one forward pass. The model is run in a custom C++ environment integrated with ROS2 Humble and an ad-hoc WiFi network for inter-node communication to communicate image embeddings $\mathbf{E}_{i}$, achieving a 15 Hz processing rate for the whole pipeline, including image pre-processing, model evaluation, communication and pose prediction.
Fig. 20 and Sec. B.4 contain additional quantitative and qualitative evaluation for the multi-robot pose-tracking experiment, consisting of two additional reference trajectories. Figure 8 static is similar to the Figure 8 dynamic trajectory elaborated in the main paper, but all robots face the same direction, resulting in two shifted Figure 8s for the two follower robots. Rectangle dynamic is also similar to the Figure 8 dynamic trajectory elaborated in the main paper, but instead of a Figure 8, the reference is a rounded rectangle, with robots moving with up to 0.8 m/s.